OpenStudy (anonymous):

Add or subtract. Enter your answer in standard form. (7u2 + 5u + 2) − (u3 + 2u2 + 1) =

OpenStudy (redneckgirl101):

−u3+5u2+5u+1

OpenStudy (kendricklamar2014):

\[7u^2+5u+2−(u^3+2u^2+1)\] Distribute the negative sign: \[=7u^2+5u+2+−1(u^3+2u^2+1)\] \[=7u^2+5u+2+−1u^3+−1(2u^2)+(−1)(1)\] \[=7u^2+5u+2+−u^3+−2u^2+−1\] Combine like terms: \[=7u^2+5u+2+−u^3+−2u^2+−1\] \[=(−u^3)+(7u^2+−2u^2)+(5u)+(2+−1)\] \[=−u^3+5u^2+5u+1\] @dsibilly

OpenStudy (anonymous):

ty

OpenStudy (kendricklamar2014):

Your welcome :)

OpenStudy (redneckgirl101):

youtr welcome

TheSmartOne (thesmartone):

Don't forget to rate the qualified helper!